# Simulation studies of the isovector reorientation effect of deuteron scattering on heavy target

### v2

@@ III.3 The data of the Analysing Power and the Simulations
In the paper Sekiguchi et al. (2017), Sekiguchi et al. provided the analyzing power data for the D(p,EL)D process, as replotted in Figure 4.
Figure 4: The relationship between the analyzing power and the center-of-mass scattering angle, data taken from Sekiguchi et al. (2017).
We will perform the experiment with the deuteron beam intensity of $1.6\times 10^{-3}$ pnA (or $10^{7}$ pps). In order to monitor the beam polarization, a $1000$ mg/cm2 thick polyethylene (CH2) target is used in the scattering chamber of the plolarimeter, which is designed to measure the counts of the elastic scattering deuteron and recoiled proton. The parameters for the detector placement are listed in Table 1. It is designed to detect the recoil protons at two positions $\theta_{1}$ and $\theta_{2}$. Figure 5 shows the coverage of the detectors and the corresponding energy.
@@
Particle $\theta_{\text{lab}}$ Distance (mm) Width $\theta$(mm) Width $\phi$(mm)
Proton $\theta_{1}$ =55.9°55.9°55.9  ° 600 20.0 20
Proton $\theta_{2}$ = 11.3°11.3°11.3  ° 600 20.0 20
Deuteron 20.87°20.87°20.87  ° 500 50 40
Figure 5: The correlation between the energy and scattering angle of the deuteron and the recoil proton. The location and the angular coverage of the detectors are represented by the colored squares.
@@
Figure 6: Ratio of the total detector counts at $\theta_{1}$ and $\theta_{2}$ during 30 min. The error bars represent the statistic uncertainty.
Figure 7: The relationship between $R_{LRUD}$ and the tensor polarization $p_{y^{\prime}y^{\prime}}$ during 30 min. The error bars represent the statistic uncertainty.
The simulation results indicate that the tensor polarization can be monitored with a precision of 0.1. This level of precision is sufficient for our experimental requirements. The monitoring method is effective and can be applied to other experiments involving tensor polarization.
@@ IV Event Generator
The scattering of polarized deuteron on heavy target is simulated by the ImQMD model, which is an extended version of QMD model suited for the simulation of heavy ion collisions as well as nucleon-induced reactions at intermediate energies. The nucleon is treated as a Gaussian wave package and follows the Hamiltonian equation of motion in the mean field. The nuclear potential energy functional including Skyrme potential energy density is written in
@@
where the isospin asymmetry is $\delta=(\rho_{\rm n}-\rho_{\rm p})/(\rho_{\rm n}+\rho_{\rm p})$, here $\rho_{\rm n}$ and $\rho_{\rm p}$ are the neutron and proton densities. All parameters in Eq. (IV) can be derived from the standard Skyrme interaction parameters. Particularly, the last term of the above equation (IV) is replaced by ${C_{\rm s,p}\over 2}\left(\textstyle{\rho\over\rho_{0}}\right)^{\gamma}\delta^{2}\rho$ to mimic the density dependence of nuclear symmetry energy as density $E_{\rm sym}(\rho)$, while the isoscale part of nuclear potential is unchanged. With this formulation, the $\gamma$ parameter can characterize the stiffness of $E_{\rm sym}(\rho)$, which is written as the sum of the kinetic term and the potential term
$E_{\rm sym}(\rho)=\frac{C_{\rm s,k}}{2}\left(\frac{\rho}{\rho_{0}}\right)^{2/3}+\frac{C_{\rm s,p}}{2}\left(\frac{\rho}{\rho_{0}}\right)^{\gamma}$ (13)
@@ V Simulation of the detection at SAMURAI
V.1 Target and Detector setup
@@
We first consider the target. The target thickness is a crucial parameter in the experiment, as it directly affects the detection efficiency and the angle straggling of the particles. A thicker target can increase the number of interactions, but it also increases the angle straggling, which can complicate the momentum reconstruction. Therefore, one seeks to find a balance between the target thickness and the detection efficiency. As Figure 8 shows, the angle straggling of the target is about 0.5 degrees for a 3 mm thick target. This means that the momentum direction of the detected particles will be affected by this angle straggling, which can lead to a significant uncertainty in the momentum reconstruction. Therefore, we need to carefully consider the target thickness and its impact on the momentum reconstruction.
Figure 8: The angle straggling of the target.
@@
Figure 9 presents the layout of the experiment installed on SAMURAI terminal. The target is placed at 3 meters to the field center on the upstream side. The neutrons are measured by the neutron wall (NEBULA), spanning an acceptance angle of ±$\pm 12^{\circ}$ with respect of the current target position. The central magnetic field strength is about 1.2 Tesla. To achieve a better acceptance angle for protons, the magnet is configured at a $30^{\circ}$, and the proton drift chamber (PDC) is placed at $42.5^{\circ}$. The path length from the field center to the incident surface of PDC is 3 meters. The two PDCs, followed by a time of flight (TOF) wall delivering the trigger signal and the TOF signal, covers the protons from the deuteron breakup emitted in the horizontal direction within ±$\pm 12^{\circ}$.
Figure 9: The layout of the detectors.
@@
In this experiment, we are interested in events where the deuteron breaks up into a neutron and a proton while the target nucleus remains intact. In order to suppress the backgound containing the fragmentation of the target, an active veto barrel (AVB) consisting of 32 plastic scintillator is placed surrounding the target rapidity. The trigger condition requires the coincidence of the NEBULA and TOF, vetoed by the AVB. The three-fold coincidence can suppressed the background greatly.
V.2 Beam dumper
@@
During the beam experiment, intense neutron background are expected in the NEBULA detector originating from the unreacted beam hitting the environmental material. Thus, to absorb the neutron background, a beam dumper is designed using water tank with dimensions of $2.5\times 2.5\times 3~{}{\rm m^{3}}$. An entrance in 34 × 38 × 120 ${\rm cm^{3}}$ is carved in the front to allow the beam to enter. It is placed at a $25^{\circ}$ and 7.5 m away from the center of the magnet.
V.3 Observation of the IVR effect in the current experimental design
In this section, the simulation results of feasibility to measure the IVR effect are presented. Using the ImQMD event generator, the $\vec{\mathrm{d}}+\mathrm{A}$ scattering at 190 MeV/u are simulated with the impact parameter ranging from 5 to 10 fm. Here A represents the target of 124Sn, 112Sn and 208Pb. The parameter $\gamma$ is varying from $\gamma=0.4$ to $0.8$, corresponding the slope parameter of $E_{\rm sym}(\rho)$ at $\rho=\rho_{0}$ varying approximately from 45 MeV to 75 MeV, respectively. Special care are taken in the response of the detector and the selection of the event.
@@ V.3 Observation of the IVR effect in the current experimental design
The detector response of the $\vec{\mathrm{d}}+\mathrm{A}$ scattering experiment is simulated using the GEANT4 toolkit. All detectors are modeled at the individual cell level, allowing precise extraction of both energy deposition and geometric position of each hit. The signals are then smeared according to the detector resolutions. Notably, although the reaction plane is predefined in the simulation, it is not known a priori in actual experiments. Therefore, the reaction plane must be reconstructed for both longitudinally and vertically polarized states, which is approximated by $\left(\vec{P}^{\mathrm{n}}+\vec{P}^{\mathrm{p}}\right)\times\hat{z}$. [^1] The reference frame is built as the Figure 3. For the vertically polarized case, we further require the reaction plane to be within ±$\pm 30^{\circ}$ of the horizontal to ensure a
@@
According to the Geant4 simulations using the SAMURAI setup, the breakup event number with $\gamma=0.6$ is 2,162,519 among 10 M deuterons with $b\leq 9$ fm, indicating a cross section $\sigma=550$ mb.
@@
Among these breakup events, 100,641 events are detected, indicating an overall detection efficiency of $\eta_{\rm det}=4.65\%$.
@@
where $d$ is the target thickness in mm, $I$ is the beam intensity in particle per second (pps) and $t$ is the beam time in hour.
@@
Now, we describe the simulations conducted to investigate the isovector reorientation effect and the selected observables used to quantify this phenomenon. Additionally, we analyze how these observables vary with the symmetry energy parameter, providing insights into the sensitivity of the observable to the underlying nuclear symmetry energy.
Figure 10 presents the distribution of the momentum of the breakup neutrons and protons in the reaction plane ( along $x$-axis) with different $\gamma$ values. The abscissa is the total momentum of the proton and neutron on the reaction plane $P_{x}^{\mathrm{n}}+P_{x}^{\mathrm{p}}$, which measures approximately the impact parameter $b$. Here $P_{x}^{\mathrm{p}}$ and $P_{x}^{\mathrm{n}}$ are the x-momentum of the proton and neutron, respectively. As one inspects the plots from the left to right, when $b<7$ fm, the breakup events is rare because the target is disintegrated. When the $b$ increases beyond 7 fm, the elastic breakup cross section increases, and the deflection of the deuteron described by $P_{x}^{\mathrm{n}}+P_{x}^{\mathrm{p}}$ decreases with $b$. The ordinate represents the momentum difference projected on the reaction plane between the proton and neutron $P_{x}^{\mathrm{p}}-P_{x}^{\mathrm{n}}$, reflecting the total effect of the isovector force and the Coulomb force acted on the neutron and the proton. With the $E_{\rm sym}(\rho)$ becomes stiffer by increasing $\gamma$ parameter, the distribution of momentum difference becomes less populated at the side of $P_{x}^{\mathrm{p}}-P_{x}^{\mathrm{n}}>0$.
@@
Finally, Fig. 13 summarizes the dependence of $R$ on the $\gamma$ parameter of polarized deuteron beam on different target. Some interesting features are visible. First, in the case of $z^{\prime}$ polarization ($p_{zz}$), it is shown that the sensitivity of $R$ on $\gamma$ is much more pronounced in 124Sn target than in 208Pb because the Coulomb effect is bigger in the latter and counteracts partially against the isospin effect. Next, the sensitivity is enhanced in the neutron rich target 124Sn in comparison to that in the less neutron rich target 112Sn. Last, the dependence of $R$ on $\gamma$ is stronger with the $z^{\prime}$ polarized deuteron beam than that with $y^{\prime}$ polarization. In all cases, the results demonstrates that the observable $R$ depends on $\gamma$ with significant sensitivity, making it a robust and sensitive probe of the isovector reorientation effect. Thus, the observable $R$ can serve as a clean probe to constrain the stiffness of the symmetry energy $E_{\mathrm{sym}}(\rho)$ in nuclear matter.
@@ VI Conclusion
The experiment of polarized deuteron scattering on various heavy target aims to investigate the isovector reorientation (IVR) effect and its inferrence of the stiffness of nuclear symmetry energy at sub-saturation density. The experiment is proposed to run at the SAMURAI spectrometer at RIKEN, utilizing a polarized deuteron beam with an energy of 190 MeV/u. In this paper, we demonstrate the method to monitor the $p_{z^{\prime}z^{\prime}}$ polarization and verify the effectiveness of $p_{y^{\prime}y^{\prime}}$ polarization monitoring proposed by previous studies. We have designed the detector setup and optimized the target position to maximize the detection efficiency for both neutrons and protons. The responses of the detectors to these particles were thoroughly analyzed, along with simulations of the background and the beam dumper to minimize the interference from background. It is found that, the IVR effect, measured by the observable $R$, is more pronounced in neutron rich target 124Sn than in neutron deficient target 112Sn, and more pronounced in 124Sn than in 208Pb. Moreover, the $z^{\prime}$-polarized beam is more favored than the $y^{\prime}$-polarized beam. In all cases, the IVR effect can be clearly observed in the simulations. The results demonstrate the feasibility of the experiment.

### v3

@@ III.3 The data of the Analysing Power and the Simulations
In the paper Sekiguchi et al. (2017), the analyzing power data for the deutron proton elastic scatering, D(p,EL)D process, is plotted in Figure 4.
Figure 4: The relationship between the analyzing power and the center-of-mass scattering angle $\theta_{\mathrm{c.m.}}$, data taken from Sekiguchi et al. (2017).
We will perform the experiment with the deuteron beam intensity of $1.6\times 10^{-3}$ pnA (or $10^{7}$ pps).Here, pnA stands for particle nano-Ampere, a unit of beam current would be measured if all beam ions were singly charged, and pps stands for particles per second. In order to monitor the beam polarization, a $1000$ mg/cm2 thick polyethylene (CH2) target is used in the scattering chamber of the polarimeter, which is designed to measure the counts of the elastic scattering deuteron and recoiled proton. The parameters for the detector placement are listed in Table 1. It is designed to detect the recoil protons at two positions $\theta_{1}$ and $\theta_{2}$. Figure 5 shows the coverage of the detectors and the corresponding energy.
@@
Particle $\theta_{\text{lab}}$ Distance (mm) Width $\theta$($m$) Width $\phi$($m$)
Proton $\theta_{1}$ =55.9° 600 20.0 20
Proton $\theta_{2}$ = 11.3° 600 20.0 20
Deuteron 20.87° 500 50 40
Figure 5: The correlation between the (a) energy and (b) scattering angle of the deuteron and the recoil proton. The location and the angular coverage of the detectors are represented by the colored squares.
@@
Figure 6: Ratio of the total detector counts at $\theta_{1}$ and $\theta_{2}$ during 30 min. The error bars (black lines inside the blue markers) represent the statistical uncertainty. Due to the large statistics collected in 30 min, these error bars are much smaller than the $\sim 10\%$ variation in tensor polarization that can be resolved.
Figure 7: The relationship between $R_{LRUD}$ and the tensor polarization $p_{y^{\prime}y^{\prime}}$ during 30 min. The error bars (black lines inside the blue markers) represent the statistical uncertainty.Thanks to the high statistics over the 30 min acquisition, these uncertainties are much smaller than the change caused by $\sim 10\%$ polarization variation.
The simulation results indicate that the tensor polarization can be monitored with a precision of 10%, which meets our experimental requirements. Importantly, the data points in Figures 6 and 7 include statistical error bars. The uncertainty at each point is substantially smaller than the typical 10% variation in polarization that we aim to resolve. This demonstrates that the statistical fluctuations would not mask a 10% change in tensor polarization. Therefore, the monitoring method is robust and can be reliably applied to similar experiments involving tensor polarization.
@@ IV Event Generator
The scattering of polarized deuteron on heavy target is simulated by the ImQMD model, which is an extended version of QMD model suited for the simulation of heavy ion collisions as well as nucleon-induced reactions at intermediate energies. The nucleon is treated as a Gaussian wave packet and follows the Hamiltonian equation of motion in the mean field. The nuclear potential energy functional including Skyrme potential energy density is written in
@@
where the isospin asymmetry is $\delta=(\rho_{\rm n}-\rho_{\rm p})/(\rho_{\rm n}+\rho_{\rm p})$, here $\rho_{\rm n}$ and $\rho_{\rm p}$ are the neutron and proton densities Ou et al. (2008). All parameters in Eq. IV can be derived from the standard Skyrme interaction parameters. Particularly, the last term of the above equation IV is replaced by ${C_{\rm s,p}\over 2}\left(\textstyle{\rho\over\rho_{0}}\right)^{\gamma}\delta^{2}\rho$ to mimic the density dependence of nuclear symmetry energy as density $E_{\rm sym}(\rho)$, while the isoscale part of nuclear potential is unchanged. With this formulation, the $\gamma$ parameter can characterize the stiffness of $E_{\rm sym}(\rho)$, which is written as the sum of the kinetic term and the potential term
$E_{\rm sym}(\rho)=\frac{C_{\rm s,k}}{2}\left(\frac{\rho}{\rho_{0}}\right)^{2/3}+\frac{C_{\rm s,p}}{2}\left(\frac{\rho}{\rho_{0}}\right)^{\gamma},$ (14)
@@ V Simulation of the detection at SAMURAI
V.1 SAMURAI Magnet
[…]
The entire spectrometer system, which also includes beam line detectors, heavy fragment detectors, and a vacuum system Shimizu et al. (2013), underwent its first commissioning in March 2012. During this commissioning period, the SAMURAI spectrometer demonstrated a rigidity resolution of approximately 1/1500 for RI beams with rigidities of up to 2.4 GeV/c, a testament to the high-performance capabilities of its magnetic component Kobayashi et al. (2013). The system’s ability to perform precise measurements is further enhanced by its detectors for both neutrons and charged particles. Projectile-rapidity nucleons, particularly neutrons, are detected using the NEBULA array.
V.2 Target and Detector setup
@@
We first consider the target. The target thickness is a crucial parameter in the experiment, as it directly affects the detection efficiency and the angular straggling of the particles. A thicker target can increase the number of interactions, but it also increases the angular straggling, which can complicate the momentum reconstruction. Therefore, one seeks to find a balance between the target thickness and the detection efficiency. As Figure 8 shows, the angular straggling is about $0.5^{\circ}$ for deutron and $1^{\circ}$ for proton for a 3 mm thick target. This means that the momentum direction of the detected particles will be affected by this angular straggling, which can lead to a significant uncertainty in the momentum reconstruction. Therefore, we need to carefully consider the target thickness and its impact on the momentum reconstruction.
Figure 8: The angular straggling of the target. The values shown are the expectation values calculated by lisecute++.
@@
Figure 9 presents the layout of the experiment installed on the SAMURAI terminal. The target is placed at 3 meters to the field center on the upstream side. Neutrons are detected by the NEBULA Nakamura and Kondo (2016) wall—a large-area neutron detector composed of stacked plastic-scintillator bars (each $\sim$12 cm $\times$ 12 cm $\times$ 180 cm) with PMTs on both ends (and preceding VETO counters)—providing time-of-flight and position information over a wide acceptance. The neutrons are measured by the neutron wall (NEBULA), spanning an acceptance angle of $\pm 12^{\circ}$ with respect to the current target position, since our target is closer to the NEBULA than the default set SAMURAI Collaboration (2005). The central magnetic field strength is about 1.2 Tesla. To achieve a better acceptance angle for protons, the magnet is configured at a $30^{\circ}$, and the proton drift chamber (PDC) is placed at $42.5^{\circ}$. The PDC consists of two downstream drift chambers using a cathode-readout method (with U, V, X cathode orientations) to provide high-resolution tracking of projectile-rapidity protons. The path length from the field center to the incident surface of PDC is 3 meters. The two PDCs cover the protons from the deuteron breakup emitted in the horizontal direction within around $\pm 12^{\circ}$ with respect to the NEBULA.
Figure 9: The layout of the particle detector. This configuration is designed to detect the neutron and proton from the deuteron breakup. The target is placed at 3 meters to the field center on the upstream side. Other detectors are shown in the Figure. 2.
@@
In this experiment, we are interested in events where the deuteron breaks up into a neutron and a proton while the target nucleus remains intact. In order to suppress the background containing the fragmentation of the target, an active veto detector (VD) consisting of 32 plastic scintillator is placed surrounding the target. The trigger condition requires the coincidence of the NEBULA and SBT, vetoed by the VD. The three-fold coincidence can suppressed the background greatly.
V.3 Beam dumper
@@
During the beam experiment, intense neutron backgrounds are expected in the NEBULA detector originating from the unreacted beam hitting the environmental material. Thus, to absorb the neutron background, a beam dumper is designed using water tank with dimensions of $2.5\times 2.5\times 3~{\rm m^{3}}$. An entrance in 34 × 38 × 120 $\rm cm^{3}$} is carved in the front to allow the beam to enter. It is placed at a $25^{\circ}$ and 7.5 m away from the center of the magnet.
V.4 Observation of the IVR effect in the current experimental design
In this section, the simulation results of the feasibility to measure the IVR effect are presented. Using the ImQMD event generator, the $\vec{\mathrm{d}}+\mathrm{A}$ scattering at 190 MeV/u are simulated with the impact parameter ranging from 5 to 10 fm. In the experimental analysis, this range of impact parameters can be selected by applying cuts on observables correlated with $b$, such as the total transverse momentum of the breakup pair, $P_{x}^{\mathrm{n}}+P_{x}^{\mathrm{p}}$, and by using centrality detectors like VD. The target ”A” is chosen for 124Sn, 112Sn and 208Pb. The parameter $\gamma$ is varying from $\gamma=0.4$ to $0.8$, as defined in Eq. 14, (The parameter $\gamma$ characterizes the stiffness of the symmetry energy as shown in Eq. 14, and $b$ is the impact parameter.)
@@ V.4 Observation of the IVR effect in the current experimental design
The detector response of the $\vec{\mathrm{d}}+\mathrm{A}$ scattering experiment is simulated using the GEANT4 toolkit. All detectors are modeled at the individual sub-detector level, allowing precise extraction of both energy deposition and geometric position of each hit. The signals are then smeared according to the detector resolutions. Notably, although the reaction plane , determined by the incident and scattered particle momenta, is predefined in the simulation, it is not known a priori in actual experiments. Therefore, the reaction plane must be reconstructed for both longitudinally and vertically polarized states, which is approximated by $\left(\vec{P}^{\mathrm{n}}+\vec{P}^{\mathrm{p}}\right)\times\hat{z}$. [^1] The reference frame is built as the Figure 3.
@@
According to the GEANT4 simulations using the SAMURAI setup, totally 2.16M breakup events with stiffness parameter $\gamma=0.6$ is recorded among 10 M deuterons with the impact parameter $b\leq 9$ fm, indicating a cross section $\sigma=550$ mb.
@@
Among these breakup events, 100K events are detected, indicating an overall detection efficiency of $\eta_{\rm det}=4.65\%$.
@@
where $d$ is the target thickness in mm, $I$ is the beam intensity in particles per second (pps) and $t$ is the beam time in hour.
@@
In the following, we describe the simulations conducted to investigate the isovector reorientation effect and the selected observables used to quantify this phenomenon. Additionally, we analyze how these observables vary with the symmetry energy parameter, providing insights into the sensitivity of the observable to the underlying nuclear symmetry energy.
Figure 10 presents the distribution of the momentum of the breakup neutrons and protons in the reaction plane ( along $x$-axis) with different $\gamma$ values. The abscissa is the total momentum of the proton and neutron on the reaction plane $P_{x}^{\mathrm{n}}+P_{x}^{\mathrm{p}}$, which measures approximately the impact parameter $b$. Here $P_{x}^{\mathrm{p}}$ and $P_{x}^{\mathrm{n}}$ are the x-momentum of the proton and neutron, respectively. As one inspects the plots from left to right, when $b<7$ fm, the breakup events is rare because the target is disintegrated. When the $b$ increases beyond 7 fm, the elastic breakup cross section increases, and the deflection of the deuteron described by $P_{x}^{\mathrm{n}}+P_{x}^{\mathrm{p}}$ decreases with $b$. The ordinate represents the momentum difference projected on the reaction plane between the proton and neutron $P_{x}^{\mathrm{p}}-P_{x}^{\mathrm{n}}$, reflecting the total effect of the isovector force and the Coulomb force acting on the neutron and the proton. With the $E_{\rm sym}(\rho)$ becomes stiffer by increasing $\gamma$ parameter, the distribution of momentum difference becomes less populated at the side of $P_{x}^{\mathrm{p}}-P_{x}^{\mathrm{n}}>0$.
@@
Finally, Fig. 13 summarizes the dependence of $R$ on the $\gamma$ parameter of the polarized deuteron beam on different target. Some interesting features are visible. First, in the case of $z^{\prime}$ polarization ($p_{zz}$), it is shown that the sensitivity of $R$ on $\gamma$ is much more pronounced in 124Sn target than in 208Pb because the Coulomb effect is bigger in the latter and counteracts partially against the isospin effect. Next, the sensitivity is enhanced in the neutron rich target 124Sn in comparison to that in the less neutron rich target 112Sn. Last, the dependence of $R$ on $\gamma$ is stronger with the $z^{\prime}$ polarized deuteron beam than that with $y^{\prime}$ polarization. In all cases, the results demonstrates that the observable $R$ depends on $\gamma$ with significant sensitivity, making it a robust and sensitive probe of the isovector reorientation effect. Thus, the observable $R$ can serve as a clean probe to constrain the stiffness of the symmetry energy $E_{\mathrm{sym}}(\rho)$ in nuclear matter.
@@ VI Conclusion
The experiment of polarized deuteron scattering on various heavy target aims to investigate the isovector reorientation (IVR) effect and its inference of the stiffness of nuclear symmetry energy at sub-saturation density. The experiment is proposed to run at the SAMURAI spectrometer at RIKEN, utilizing a polarized deuteron beam with an energy of 190 MeV/u. In this paper, we demonstrate the method to monitor the $p_{z^{\prime}z^{\prime}}$ polarization and verify the effectiveness of $p_{y^{\prime}y^{\prime}}$ polarization monitoring proposed by previous studies. We have designed the detector setup and optimized the target position to maximize the detection efficiency for both neutrons and protons. The responses of the detectors to these particles were thoroughly analyzed, along with simulations of the background and the beam dumper to minimize the interference from the background. It is found that, the IVR effect, measured by the observable $R$, is more pronounced in the neutron-rich target 124Sn than in the neutron-deficient 112Sn, and also more pronounced in 124Sn than in 208Pb. Moreover, the $z^{\prime}$-polarized beam is more favored than the $y^{\prime}$-polarized beam. In all cases, the IVR effect can be clearly observed in the simulations. The simulation predictions demonstrate the feasibility of the experiment.